OpenStudy (anonymous):

Please help: Solve: 4x^2 + 8x - 5 = 0

OpenStudy (anonymous):

@mariah95 Here you go

OpenStudy (anonymous):

okay got you give me a sec..

OpenStudy (anonymous):

why did put a ^ thing

OpenStudy (anonymous):

To say it was a square.

OpenStudy (anonymous):

oh okay

OpenStudy (radar):

Complete the square method. 1. Divide thru by 4: x^2 +2x -5/4 = 0 2. Add 5/4 to both sides getting: x^2 + 2x = 5/4 3. Take 1/2 of the x coefficient and square it adding it to both sides. Getting: x^2 + 2x + 1 = 5/4 + 1 Simplify: (x + 1)^2 = 9/4 Take square root of both sides: x + 1 = =/- 3/2 solve for x, x + 1 = -3/2 x=-1 - 3/2 = -5/2 (1st root) x + 1 = 3/2 x=-1 + 3/2 x= 1/2 2nd root

OpenStudy (anonymous):

Sir how would that be if u solved it as a Quadratic

OpenStudy (radar):

Should get the same two roots. I just liked that way better. Does the problem require you to use quadratic? If it does, please do so.

OpenStudy (anonymous):

I Believe i can do it that way but my teachers have not taught it to use that way

OpenStudy (radar):